OpenStudy (anonymous):

simplify the rational expression... x^2-64/8+x

OpenStudy (anonymous):

x-8

OpenStudy (anonymous):

\[\frac{x^2-64}{x+8}=\frac{(x+8)(x-8)}{x+8}=x-8\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw